# Attractor Memory for Long-Term Time Series Forecasting: A Chaos Perspective

### v4

@@ A.1 Taken Theorem
That is, there is a diffeomorphism $\varphi$ that maps $\mathcal{A}$ into $\mathbb{R}^{N}$ such that the derivative of $\varphi$ has full rank. A delay embedding theorem uses an observation function to construct the embedding function. An observation function $h:\mathcal{M}\rightarrow\mathbb{R}$ must be twice-differentiable and associate a real number to any point of the attractor $\mathcal{A}$. It must also be typical, so its derivative is of full rank and has no special symmetries in its components. The delay embedding theorem states that the function
@@
is an embedding of the strange attractor $\mathcal{A}$ in $\mathbb{R}^{N}$.
Suppose the $d$-dimensional state vector $x_{t}$ evolves according to an unknown but continuous and (crucially) deterministic dynamic. Suppose, too, that the one-dimensional observable $y$ is a smooth function of $x$, and "coupled" to all the components of $x$. Now at any time we can look not just at the present measurement $y(t)$, but also at observations made at times removed from us by multiples of some lag $\tau:y_{t+\tau},y_{t+2\tau}$, etc. If we use $m$ lags, we have a $m$-dimensional vector. One might expect that, as the number of lags is increased, the motion in the lagged space will become more and more predictable, and perhaps in the limit $m\rightarrow\infty$ would become deterministic. In fact, the dynamics of the lagged vectors become deterministic at a finite dimension; not only that, but the deterministic dynamics are completely equivalent to those of the original state space (precisely, they are related by a smooth, invertible change of coordinates, or diffeomorphism). In fact, the theorem says that determinism appears once you reach dimension $12d+$ 1, and the minimal embedding dimension is often less.
@@ A.3 Orthogonal Polynomials
A.4 Piecewise Projection
and back-projection can be achieved through the Moore-Penrose inverse $H^{\dagger}$. By taking the inner product with $u$ on both sides in Eq. 20, we can project the state representation $x(t)$ between $\mathcal{G}_{\theta}^{k}$ space and the $\mathcal{G}_{2\theta}^{k}$ space.
@@ Appendix B Proof
$\bm{A}=\mathsf{diag}\{-1,-1,\dots\}$ is a rough approximation of shifted Hippo-LegT [11] matrix.
@@
Hippo 3 provides a mathematical framework for deriving the $\bm{AB}$ matrix for polynomial projection. The mainstream SSMs [10] typically initialize matrix $\bm{A}=\mathsf{diag}\{-1,-2,-3,-4,\dots\}$, representing the negative real diagonal elements of the normalized Hippo-LegS matrix (21). Since the LegS matrix is a mathematical approximation of exponentially decaying Legendre polynomials, initializing $\bm{A}=\mathsf{diag}\{-1,-2,-3,-4,\dots\}$ can be seen as a rough approximation of exponential decay. Similarly, for the normalized Hippo-LegT matrix (22), which approximates a uniform measure, we can consider $\bm{A}=\mathsf{diag}\{-1,-1,\dots\}$ as a rough approximation of a finite window of Legendre polynomials. The use of negative values for the elements is to ensure gradient stability during training.
@@ C.1 Phase Space Reconstruction
Phase space reconstruction is a crucial technique in the analysis of dynamic systems, particularly in the study of time series data. This method transforms a one-dimensional time series into a multidimensional phase space, revealing the underlying dynamics of the system. The cross-correlation mutual information (CC mutual information) method is a statistical tool used to analyze the dependencies between two different time series. We provide a detailed mathematical description of these methods.
@@
Takens’ Embedding Theorem allows the reconstruction of a dynamic system’s phase space from a sequence of observations. The theorem states that under generic conditions, a map:
@@ D.3 Experiments Setting
All experiments are conducted on the NVIDIA RTX3090-24G and A6000-48G GPUs. The Adam optimizer is chosen. A grid search is performed to determine the optimal hyperparameters, including the learning rate from {0.0001, 0.0005, 0.001}, polynomial order from {32, 64, 256}, patch length from {8, 16}, projection level based on $logL$ and no more than 3, and primary modes from {32, 64}. The input length is 96. Due to the sensitivity of the CC method to numerical calculations, we take the average result from three iterations of the calculation.
@@ E.2 The Chaotic Nature of Datasets
Lyapunov exponents are a set of quantities that characterize the rate of separation of infinitesimally close trajectories in a dynamic system. They play a crucial role in understanding the stability and chaotic behavior of the system [21]. The formal definition of the Lyapunov exponent for a trajectory of a dynamic system is given by:
@@ E.3 Simulation for Lorenz96 (Case Study)
The Lorenz ’63 model, introduced by Edward Lorenz in 1963, is a simplified mathematical model for atmospheric convection. The model is a system of three ordinary differential equations now known as the Lorenz equations:
@@
To simulate time series generated from an unknown chaotic system, we first construct a 40-dimensional Lorenz96 system to represent the underlying chaotic system. We then use a randomly initialized linear neural network to simulate the observation function $h$. Through $h$, we map the Lorenz96-40d system, which resides in the manifold space, to a 3-dimensional Euclidean space to obtain a multivariate time series dataset with three variables. As shown in the middle of Figure 6, in the observation domain, the dynamic system structure of Lorenz96-40d is difficult to discern specific shapes, so we need to reconstruct it to 40 dimensions using the Phase Space Reconstruction (PSR) technique to study its dynamic characteristics in a topologically equivalent structure. It is worth noting that on the right side of Figure 6, we visualize the last variable of the system and find striking similarities with real-world time series, even exhibiting some periodic behaviors. This further supports our hypothesis that real-world time series are generated by underlying chaotic systems.
@@ E.4 Chaotic Modulation
In LTSF tasks, it is commonly observed that predictive performance deteriorates as the length of the forecasting window increases. This phenomenon bears a striking resemblance to the inherent challenge in long-term predictions of chaotic dynamic systems, which are highly sensitive to initial conditions. Recent studies [15] in the field of chaotic dynamic systems have highlighted that, to address the issue of gradient divergence caused by positive maximal Lyapunov exponents indicative of chaos, the implementation of teaching forcing as a method to incrementally constrain the trajectory of the dynamic system presents a straightforward yet effective framework. To enhance Hippo-RNN, we have implemented the following modifications: $\tilde{\bm{z}}_{t}:=(1-\alpha)\bm{z}_{t}+\alpha\overline{\bm{z}}_{t},~{}\bm{z}_{t}=\bm{F}_{\bm{\theta}}\left(\tilde{\bm{z}}_{t-1}\right),~{}\text{with}~{}0\leq\alpha\leq 1$, where we intervene the evolution state $\bm{z}$ by utilizing the ground truth hidden state $\overline{\bm{z}}$. As shown in Figure 7, this modification leads to notable improvements in both mainstream LTSF datasets and real-world chaotic datasets. However, the autoregressive nature of the teaching forcing methods introduces additional computational overhead and may potentially reduce its generalization capabilities, which presents a challenge for integrating it into sequence-mapping models.
E.5 Chaotic Reconstruction
As illustrated in Figure 8, we visualize the phase space forecasting results of Attraos on the Lorenz96 system. It can be observed that: Although some minor details may be missing due to the sparsity introduced by the frequency domain evolution, Attraos successfully reconstructs the chaotic dynamic structure of Lorenz96. Moreover, modeling time series based on the dynamics structure of the phase space can be viewed as a form of data augmentation, e.g., two-dimensional time figuring [42] or seasonal decomposition [45].
Figure 8: Chaotic reconstruction process for Lorenz96-40d. Left: Ground truth. Right: Prediction.
E.6 Chaos Representation (Hyper-parameter Analysis)
We validated the impact of polynomial dimensions on the model performance and observed that the chunking operation significantly reduces the reliance on polynomial space dimensions, thereby greatly reducing the number of model parameters.
Figure 9: Hyper-parameter analysis, the X-axis is polynomial order. Metric is MSE in ETTm2 dataset.
E.7 Full Results

### v7

@@ A.1 Takens Theorem
This implies that there is a diffeomorphism $\varphi$ mapping the attractor $\mathcal{A}$ into $\mathbb{R}^{N}$, such that the Jacobian matrix of $\varphi$ is of full rank. In the delay embedding process, the embedding function is constructed using an observation function. This observation function $h:\mathcal{M}\rightarrow\mathbb{R}$ needs to be twice continuously differentiable and must assign a real number to each point on the attractor $\mathcal{A}$, ensuring that it is typical, meaning its derivative is of full rank without exhibiting any special symmetries.
The delay embedding theorem states that the mapping
@@
constitutes an embedding of the strange attractor $\mathcal{A}$ into $\mathbb{R}^{N}$.
Now, consider a $d$-dimensional state vector $x_{t}$, which evolves according to an unknown but deterministic and continuous dynamic process. Suppose a one-dimensional observable $y$ exists, which is a smooth function of $x$ and is coupled to all the components of $x$. At any given time, we can observe not only the current measurement $y(t)$ but also measurements from past times separated by a lag $\tau$: $y_{t+\tau},y_{t+2\tau}$, and so on. Using $m$ such lags results in an $m$-dimensional vector. As the number of lags increases, the motion in the reconstructed space becomes more predictable, and in the limit as $m\rightarrow\infty$, the dynamics could become deterministic. In reality, the deterministic nature of the dynamics is achieved at a finite dimension, with the reconstructed dynamics being equivalent to the original system’s dynamics, related by a smooth, invertible change of coordinates (a diffeomorphism). The theorem specifically asserts that deterministic behavior emerges once the dimension reaches $12d+$ 1, with the minimal embedding dimension often being lower.
@@ Appendix B Proof
[18] $\bm{A}=\mathsf{diag}\{-1,-1,\dots\}$ is a rough approximation of shifted Hippo-LegT [11] matrix.
@@
Hippo 3 provides a mathematical framework for deriving the $\bm{AB}$ matrix for polynomial projection. The mainstream SSMs [10] typically initialize matrix $\bm{A}=\mathsf{diag}\{-1,-2,-3,-4,\dots\}$, representing the negative real diagonal elements of the normalized Hippo-LegS matrix (21). Since the LegS matrix is a mathematical approximation of exponentially decaying Legendre polynomials, initializing $\bm{A}=\mathsf{diag}\{-1,-2,-3,-4,\dots\}$ can be seen as a rough approximation of exponential decay. Similarly, for the normalized Hippo-LegT matrix (22), which approximates a uniform measure, we can consider $\bm{A}=\mathsf{diag}\{-1,-1,\dots\}$ as a rough approximation of a finite window of Legendre polynomials, which better for non-stationary time series as well [19]. The use of negative values for the elements is to ensure gradient stability during training.
@@ C.1 Phase Space Reconstruction
Phase space reconstruction is a crucial technique in the analysis of dynamical systems, particularly in the study of time series data. This method transforms a one-dimensional time series into a multidimensional phase space, revealing the underlying dynamics of the system. The cross-correlation mutual information (CC mutual information) method is a statistical tool used to analyze the dependencies between two different time series. We provide a detailed mathematical description of these methods.
@@
Takens’ Embedding Theorem allows the reconstruction of a dynamical system’s phase space from a sequence of observations. The theorem states that under generic conditions, a map:
@@ D.3 Experiments Setting
All experiments are conducted on the NVIDIA RTX3090-24G and A6000-48G GPUs. The Adam optimizer is chosen. A grid search is performed to determine the optimal hyperparameters, including the learning rate from {0.0001, 0.0005, 0.001}, model layer from {1, 2} (typically 1), polynomial order from {32, 64, 256}, patch length from {8, 16}, projection level based on $logL$ and no more than 3, primary modes is 16, and the input length is 96. Due to the sensitivity of the CC method to numerical calculations, we take the average result from three iterations of the calculation.
@@ E.2 The Chaotic Nature of Datasets
Lyapunov exponents are a set of quantities that characterize the rate of separation of infinitesimally close trajectories in a dynamical system. They play a crucial role in understanding the stability and chaotic behavior of the system [24]. The formal definition of the Lyapunov exponent for a trajectory of a dynamical system is given by:
@@ E.3 Simulation for Lorenz96 (Case Study)
The Lorenz63 system, introduced by Edward Lorenz in 1963, is a simplified mathematical model for atmospheric convection. The model is a system of three ordinary differential equations now known as the Lorenz equations:
@@
To simulate time series generated from an unknown chaotic system, we first construct a 40-dimensional Lorenz96 system to represent the underlying chaotic system. We then use a randomly initialized linear neural network to simulate the observation function $h$. Through $h$, we map the Lorenz96-40d system, which resides in the manifold space, to a 3-dimensional Euclidean space to obtain a multivariate time series dataset with three variables. As shown in the middle of Figure 7, in the observation domain, the dynamical system structure of Lorenz96-40d is difficult to discern specific shapes, so we need to reconstruct it to 40 dimensions using the Phase Space Reconstruction (PSR) technique to study its dynamic characteristics in a topologically equivalent structure. It is worth noting that on the right side of Figure 7, we visualize the last variable of the system and find striking similarities with real-world time series, even exhibiting some periodic behaviors. This further supports our hypothesis that real-world time series are generated by underlying chaotic systems.
@@ E.4 Chaotic Modulation
In LTSF tasks, it is commonly observed that predictive performance deteriorates as the length of the forecasting window increases. This phenomenon bears a striking resemblance to the inherent challenge in long-term predictions of chaotic dynamical systems, which are highly sensitive to initial conditions. Recent studies [15] in the field of chaotic dynamical systems have highlighted that, to address the issue of gradient divergence caused by positive maximal Lyapunov exponents indicative of chaos, the implementation of teaching forcing as a method to incrementally constrain the trajectory of the dynamical system presents a straightforward yet effective framework. To enhance Hippo-RNN, we have implemented the following modifications: $\tilde{\bm{z}}_{t}:=(1-\alpha)\bm{z}_{t}+\alpha\overline{\bm{z}}_{t},~{}\bm{z}_{t}=\bm{F}_{\bm{\theta}}\left(\tilde{\bm{z}}_{t-1}\right),~{}\text{with}~{}0\leq\alpha\leq 1$, where we intervene the evolution state $\bm{z}$ by utilizing the ground truth hidden state $\overline{\bm{z}}$. As shown in Figure 8, this modification leads to notable improvements in both mainstream LTSF datasets and real-world chaotic datasets. However, the autoregressive nature of the teaching forcing methods introduces additional computational overhead and may potentially reduce its generalization capabilities, which presents a challenge for integrating it into sequence-mapping models.
E.5 Full Results